Jisc Major Incidents
            
            @jiscmi
Followers
                3K
              Following
                0
              Media
                0
              Statuses
                434
              This account broadcasts information on MIs affecting the Janet network. Full details will be available at https://t.co/NlGGEExlc9 or 0300 300 2212.
              
              Harwell, Oxford
            
            
              
              Joined January 2015
            
            
           As of approximately 8:30 last night, all sites are back up and running as normal. 
          
                
                0
              
              
                
                0
              
              
                
                1
              
             A number of Jisc sites have been affected by the Microsoft Azure Front Door outage including Hedd and Chest. We are monitoring the situation closely and will provide an update at 9:00 tomorrow morning if not before. 
          
                
                1
              
              
                
                2
              
              
                
                4
              
             186395 - We have been made aware that splicing work is now complete. We are awaiting confirmation as to whether this is a permanent or temporary fix. 
           186395 - Connectivity to all services has now been restored. Repair work is not yet complete, listed services will continue to be considered at-risk until completion. Updates to follow. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186395 - Fibre repair work is ongoing to restore connectivity to all services. Updates to follow. 
           186395 - Connectivity restored to the majority of affected services at 22:22 BST, following the reboot of an optical node. The remaining services are expected to restore once fibre splicing has been completed. Updates to follow. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186395 - Connectivity restored to the majority of affected services at 22:22 BST, following the reboot of an optical node. The remaining services are expected to restore once fibre splicing has been completed. Updates to follow. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186395 - Our provider has advised fibre engineers have started splicing the 96 Fibre, services will restore as the splicing is completed at both joints within the new pits. Updates to follow. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             TT186395 - Our provider has advised sub work is being installed to the new pit locations for the fibre to be terminated within the pits. Updates to follow. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186395 - Our provider has advised Civils are preparing the ducts for the cable to be installed then fibre splicing can commence. Updates to follow. 
          
                
                0
              
              
                
                0
              
              
                
                1
              
             186395 - Our provider has confirmed that engineers are onsite and have started pulling through fibres. Due to the extent of the damage an accurate ETR is not available at this time. Updates to follow 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186395 - Our provider has given us an ETA of 18:30 for civil works and splicing team to arrive at the break location to begin repair works on the damaged fibre. Updates to follow. 
           186395 - Our provider has now confirmed plans to replace the damaged fibre with a new 96 fibre cable. We do not currently have an ETR but will continue to press for progress updates during the repair. Updates to follow. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186395 - Our provider has now confirmed plans to replace the damaged fibre with a new 96 fibre cable. We do not currently have an ETR but will continue to press for progress updates during the repair. Updates to follow. 
           186395 - Declared Bronze Major Incident We currently have an outage affecting South Region since 10:27 BST due to a fibre break caused by a telephone pole install 10km outside Wymering. We are working with our circuit provider for a resolution. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186395 - Declared Bronze Major Incident We currently have an outage affecting South Region since 10:27 BST due to a fibre break caused by a telephone pole install 10km outside Wymering. We are working with our circuit provider for a resolution. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186170 - Connectivity has resorted to all affected services connected via Erdington C-PoP at 02:16 UTC Tuesday 26th March 2024. Our engineers have checked our equipment within Erdington C-PoP and confirmed power has restored. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186170 - We have been advised by our third-party provider that the outage affecting Erdington C-PoP is related to power works to replace air circuit breakers. We expect power to be restored to Erdington C-PoP by 02:00 UTC on Tuesday, March 26th, 2024. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186170 - Loss of connectivity affecting Members' services connected via Erdington C-PoP, Our engineers are currently investigating a reason for this outage. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186152 - Engineer has gained access to PoP location 23:36 UTC. Swap of faulty part will begin imminently. Updates to follow. 
           186152 - The engineer has arrived on site, though there is a lack of external lighting to gain access to the building. The engineer has temporarily left the premises to collect a torch. 
          
                
                0
              
              
                
                0
              
              
                
                0
              
             186152 - The engineer has arrived on site, though there is a lack of external lighting to gain access to the building. The engineer has temporarily left the premises to collect a torch. 
          
                
                0
              
              
                
                1
              
              
                
                12